OpenStudy (anonymous):

Solve. 6x^2 – 24x = 126. x = –7 and x = –3 x = –7 and x = 3 x = 7 and x = –3 x = 7 and x = 3

OpenStudy (anonymous):

x = 7 and x = –3

OpenStudy (anonymous):

divide the equation by 6 x^2 - 4x = 21 x^2 - 4x - 21 = 0 (x-7)(x+3) x= 7 and x = -3